OpenStudy (anonymous):

Multiply and simplify: x-6/x^2+4x-32*x-4/x+2

OpenStudy (campbell_st):

factorise the quadratic x^2 + 4x - 32 = (x +8)(x - 4) the problem is now \[((x-6)(x-4))/(((x+8)((x-4)(x+2) \] cancel common factors \[(x - 6)/((x+8)(x+2) = (x - 6)/(x^2 +10x + 16)\]

OpenStudy (anonymous):

here are the choices to chose from: a.x-6/x+2 b.x-6/(x+8)(x+2) c.x-3/x+8

OpenStudy (anonymous):

b